The perimeter of a rectangle is 68 cm. If the width is increased by 2cm. and twice th length is decreased by 10cm, the new perimeter is 100 cm. What are the dimensions of the original rectangle?
Answer by mananth(16949)

let width be x
length =y
2(x+y)=68
divide by 2 both sides
x+y=34-------------------1
..
2*((x+2)+(2y+10)=100
divide by 2
x+2+2y+10=50
x+2y+12=50
x+2y=50-12
x+2y=38-----------------2
Subtract equation 1 from 2
x+2y-x-y=38-34
y=4..
x+y=34
plug the value of y
x+4=34
x=30
The dimensions of original rectangle are 30cm by 4cm